### v3

Abstract
In this paper we propose a novel modification of Contrastive Language-Image Pre-Training (CLIP) guidance for the task of backlit image enhancement. Our work builds on the state-of-the-art CLIP-LIT approach, which learns a prompt pair by constraining the text-image similarity between a prompt (negative/positive sample) and a corresponding image (backlit image/well-lit image) in the CLIP embedding space. Learned prompts then guide an image enhancement network. Based on the CLIP-LIT framework, we propose two novel methods for CLIP guidance. First, we show that instead of tuning prompts in the space of text embeddings, it is possible to directly tune their embeddings in the latent space without any loss in quality. This accelerates training and potentially enables the use of additional encoders that do not have a text encoder. Second, we propose a novel approach that does not require any prompt tuning. Instead, based on CLIP embeddings of backlit and well-lit images from training data, we compute the residual vector in the embedding space as a simple difference between the mean embeddings of the well-lit and backlit images. This vector then guides the enhancement network during training, pushing a backlit image towards the space of well-lit images. This approach further dramatically reduces training time, stabilizes training and produces high quality enhanced images without artifacts. Additionally, we show that residual vectors can be interpreted, revealing biases in training data, and thereby enabling potential bias correction. Code is available at https://github.com/Atmyre/RAVE
Keywords:
Backlit image enhancement Vision-language models Residual vector embedding
Accepted to ECCV 2024
1 Introduction
Figure 1: Visual comparison of results obtained by RAVE vs CLIP-LIT. RAVE produces well-lit images with fewer artifacts (note the dark green color of underexposed regions in CLIP-LIT), as shown in the zoomed-in sections. Further comparisons can be found in supplementary material.
Backlit image enhancement aims to improve images that suffer from backlighting issues. Backlighting is a phenomenon when the light source is situated behind the photographed object, often resulting in a loss of detail and contrast of some areas due to underexposure, diminishing the overall visual quality of the image. The goal of backlit image enhancement is to correct these exposure discrepancies and bring the details in the darkened areas out without overexposing the well-exposed parts of the image. Backlit image enhancement significantly improves photograph quality in uncontrolled lighting conditions. It is also vital for fields requiring clear image visibility, such as surveillance, security, and scientific imaging applications.
However, correcting backlit images is not an easy task. Manual correction requires skill using photo enhancement software, and often substantial time and effort. Automated solutions are preferred, but also face challenges. Global adjustment of the brightness level is not sufficient, as this typically overexposes regions of the image which were well lit prior to enhancement [11]. Spatially adaptive methods have appeared in the literature including early methods that rely on assumptions such as average luminance and a gray world model [8, 12] or Retinex theory [10]. More recent methods approximate well-lit image distributions via end-to-end training of deep networks [26, 3]. Despite progress in the field, considerable room for improvement exists for high quality automated backlit image enhancement.
Another challenge to high quality automatic backlit image enhancement is the lack of paired data, i.e. backlit images each with a corresponding well-lit image. It is complicated to collect such data, so approaches that work with unpaired data are imperative. Recently, CLIP-LIT[14] was proposed, which successfully utilizes CLIP[19] model guidance for training a backlit enhancement model with unpaired data. CLIP-LIT uses prompt learning techniques to provide CLIP guidance. More specifically, it constructs two learnable text prompts, which are trained to have CLIP embeddings close to the well-lit and backlit images, respectively. These learnt prompts then guide the image enhancement model during training. Next, several iterations of additional prompt correction followed by enhancement model fine-tuning using the updated prompts are required to produce the final image enhancement model. This approach is the current state-of-the-art in backlit image enhancement with training on unpaired data.
In this work, we demonstrate that prompt training is not the most efficient way to implement the CLIP guidance, and propose two novel methods, named CLIP-LIT-Latent and ResiduAl Vector Embedding (RAVE). In CLIP-LIT-Latent, we train vectors corresponding to well-lit and backlit images directly in the CLIP latent space rather than in text embedding space. We show that this approach has similar or superior performance when trained on paired or unpaired image data. Further, in RAVE, we form the guidance vector by subtracting the mean of CLIP latent vectors of backlit images from the mean of CLIP latent vectors of well-lit images. This residual vector points in a direction of moving backlit images to well-lit images in the CLIP embedding space. This vector is then used as guidance for the image enhancement model during training. We show that such guidance enables the enhancement model to produce high quality images, while requiring much less training time. While having all these benefits, both proposed CLIP-LIT-Latent and RAVE retain all the advantages of the original CLIP-LIT method, being lightweight for inference and suitable for training on paired or unpaired data. Moreover, we show that residual vector used for guidance in RAVE is interpretable, which opens up new possibilities for overcoming biases in the training data.
The contributions of this paper can be summarized as follows:
  1. 1.
    Inspired by CLIP-LIT, we present two novel approaches to CLIP guidance (CLIP-LIT-Latent and RAVE) working directly in the latent space for the task of backlit image enhancement.
  2. 2.
    We show that these approaches result in similar or better quality according to quantitative metrics in both settings of training on paired or unpaired data, and superior visual quality of resulting images, while requiring considerably less time to train.
  3. 3.
    We demonstrate that the embedding used by RAVE for guidance is interpretable, and its interpretation can reveal biases in the training data.
2 Related work
[…]
Backlit Image Enhancement. For the task of backlit image enhancement, many approaches have been proposed. Some techniques aim to identify backlit regions and correct them separately from the rest of the image[13]. Methods based on the Retinex[10] theory divide an image into a reflectance image and an illumination map, and both parts are then enhanced and then combined back together to form the resulting enhanced image. Another approach to enhancing backlit images is using end-to-end deep learning approaches, as in Restormer[26] or Retinexformer[3] methods. Combined techniques based on Retinex theory and deep learning were proposed for both backlit and low-light image enhancement, such as Diff-Retinex[25]. These methods decomposing an input image into a reflectance image and an illumination map and apply diffusion methods for enhancement before combining them to produce the output image. Finally, another set of methods are based on high dynamic range (HDR) imaging which combine multiple exposures to produce a composite well-lit image[20, 28, 4]. However, in this work we only focus on single-image methods where back-lit correction is applied as a post-process.
Fundamentally, backlit image correction is ill-posed, in that for a single backlit image, multiple well-lit variants with high perceptual quality exist. Relatedly, generation of paired training data requires hiring professional photographers, which can be expensive. To help address this problem, recently the BAID[15] dataset was proposed, which contains paired images generated by professional photographers. Nevertheless, development of methods for backlit image enhancement, which can be trained using non-paired data remains essential.
[…]
3 Methodology
Our work is built on top of the CLIP-LIT[14] approach, which uses CLIP guidance for training a backlit image enhancement model. We propose two modifications of CLIP guidance, which reduce training time, stabilize training and produce high quality enhanced images with fewer artifacts compared to CLIP-LIT.
We start in Section 3.1 by reviewing the architecture of original CLIP-LIT method on which our work is based. Then, in Sections 3.2 and 3.3 we introduce our proposed methods CLIP-LIT-Latent and RAVE.
3.1 CLIP-LIT overview
(a) Initial Prompts and Enhancement Training
(b) Prompt Refinement and Enhancement Tuning
Figure 2: Overview of the original CLIP-LIT approach and proposed CLIP-LIT-Latent. (a) depicts the first stage of training, which consists of prompt or latent vector initialization and the initial training of an enhancement network. (b) shows the second stage, where prompt/latent vector refinement and enhancement model fine-tuning are iteratively repeated. Blue and red boxes are related to CLIP-LIT and CLIP-LIT-Latent.
The CLIP-LIT framework (Fig.  2) consists of two stages, described below.
Initial Prompts and Enhancement Training
In this stage, two prompts $T_{p},T_{n}\in\mathbb{R}^{N\times 512}$ representing positive and negative prompts are randomly initialized, where $N$ is a number of tokens in each prompt. Then they are tuned so that the CLIP embedding of $T_{p}$ is close in terms of the dot product to the CLIP embeddings of well-lit training images, and CLIP embedding of $T_{n}$ is close to the CLIP embeddings of backlit images from the training data. The binary cross-entropy loss of classifying the backlit and well-lit images based on $T_{p}$ and $T_{n}$ is used for this purpose:
$\begin{split}L_{\text{initial}}=-(y\log(\widehat{y})+(1-y)\log(1-\widehat{y}))\\ \widehat{y}=\frac{e^{\cos(\Phi_{\text{image}}(I),\Phi_{\text{text}}(T_{p}))}}{\sum_{i\in\{n,p\}}e^{\cos(\Phi_{\text{image}}(I),\Phi_{\text{text}}(T_{i}))}}\end{split}$ (1)
Here $I$ is a well-lit or backlit input image with label $y\in\{0,1\}$ with 0 corresponding to backlit and 1 corresponding to well-lit; $\Phi_{\text{image}}$ is a CLIP[19] image encoder, $\Phi_{\text{text}}$ is a CLIP text encoder; $\cos$ is a cosine similarity function.
Using these initialized prompts $T_{p}$ and $T_{n}$ it is now possible to train an image enhancement network with a CLIP-guided loss. The enhancement model is a Unet[21] which takes backlit image $I_{b}$ as input and outputs an estimated illumination map $I_{i}\in\mathbb{R}^{H\times W\times 1}$. The final enhanced image $I_{t}$ is then obtained as $I_{t}=I_{b}/I_{i}$. The image enhancement network is trained by combining two losses:
$L_{\text{enhance}}=L_{\text{clip}}+\omega L_{\text{identity}}$ (2)
Here $L_{\text{clip}}$ measures the similarity between the enhanced image and the learned prompts $T_{p}$ and $T_{n}$ in the CLIP latent space. $L_{\text{identity}}$ is an identity loss designed to ensure that the enhanced image is similar to the original input image in terms of content and structure. The parameter $\omega$ balances the two losses. These losses are defined as:
$L_{\text{clip}}=\frac{e^{\cos(\Phi_{\text{image}}(I_{t}),\Phi_{\text{text}}(T_{n}))}}{\sum_{i\in\{n,p\}}e^{\cos(\Phi_{\text{image}}(I_{t}),\Phi_{\text{text}}(T_{i}))}}$ (3)
$L_{\text{identity}}=\sum_{l=0}^{k}\alpha_{l}\cdot\|\Phi^{l}_{\text{image}}(I_{b})-\Phi^{l}_{\text{image}}(I_{t})\|_{2}$ (4)
where $\Phi^{l}$ is $l^{th}$ layer of the CLIP image encoder, and $\alpha_{l}$ is its corresponding weight. $k$ is number of layers used to calculate the metric.
Prompt Refinement and Enhancement Tuning
CLIP-LIT applies a second stage where several cycles of prompt refinement and enhancement network tuning are performed to improve the accuracy of the learned prompts.
First, learned prompts $T_{p}$ and $T_{n}$ are fine-tuned to provide better guidance for fine-tuning the image enhancement model. A margin ranking loss is used to update the prompts based on images given by the current enhancement model:
$\begin{split}L^{1}_{\text{prompt}}&=\max(0,S(I_{w})-S(I_{b})+m_{0})+\max(0,S(I_{t-1})-S(I_{b})+m_{0})\\ &+\max(0,S(I_{w})-S(I_{t})+m_{1})+\max(0,S(I_{t})-S(I_{t-1})+m_{2})\end{split}$ (5)
where $S(I)$ is a negative similarity score between the prompt pair and an image defined as:
$S(I)=\frac{e^{\cos(\Phi_{\text{image}}(I),\Phi_{\text{text}}(T_{n}))}}{\sum_{i\in\{n,p\}}e^{\cos(\Phi_{\text{image}}(I),\Phi_{\text{text}}(T_{i}))}}$ (6)
Here $m_{0}\in[0,1]$ is the margin between the score of well-lit or enhanced result and the backlit image in the CLIP embedding space; $m_{1}\in[0,1]$ is the margin between the score of the enhanced result and the well-lit image in the CLIP embedding space; $m_{2}\in[0,1]$ is the margin between the newly enhanced result $I_{t}$ and previously enhanced result $I_{t-1}$.
After that, a stage of enhancement model fine-tuning is performed. This follows the same process as initial enhancement model training described above. The process of prompt refinement and enhancement tuning is repeated several times until satisfactory results are obtained.
3.2 CLIP-LIT-Latent
Like CLIP-LIT, CLIP-LIT-Latent also consists of two stages of initial prompts and enhancement training and prompt refinement and enhancement tuning. Fig.  2 illustrates the CLIP-LIT-Latent framework.
Instead of learning prompts in the text embedding space, CLIP-LIT-Latent learns a pair of positive/negative vectors directly in the CLIP latent space. In the prompt initialization phase, we randomly initialize two vectors $E_{p},E_{n}\in\mathbb{R}^{512}$. These vectors are then tuned so that $E_{p}$ is close in terms of the dot product to the CLIP embedding of well-lit images from the training data, and $E_{n}$ is close to the CLIP embedding of backlit images. As in original CLIP-LIT approach, binary cross-entropy loss for classifying the backlit and well-lit images based on $E_{p}$ and $E_{n}$ is used for this purpose:
$\begin{split}L_{\text{initial}}=-(y\log(\widehat{y})+(1-y)\log(1-\widehat{y}))\\ \widehat{y}=\frac{e^{\cos(\Phi_{\text{image}}(I),E_{p})}}{\sum_{i\in\{n,p\}}e^{\cos(\Phi_{\text{image}}(I),E_{i})}}\end{split}$ (7)
where $I$ is a well-lit or backlit input image with label $y\in{0,1}$ with 0 corresponding to backlit and 1 corresponding to well-lit; $\Phi_{\text{image}}$ is a CLIP image encoder.
Based on these initialized vectors $E_{p}$ and $E_{n}$, an enhancement network is trained with a CLIP-guided loss. As before, the enhancement model is a Unet which takes backlit image $I_{b}$ as input and outputs an estimated illumination map $I_{i}\in\mathbb{R}^{H\times W\times 1}$. And, as before, it is trained using $L_{\text{enhance}}$ loss, where the $L_{\text{clip}}$ component is changed as follows:
$L_{\text{clip}}=\frac{e^{\cos(\Phi_{\text{image}}(I_{r}),E_{n})}}{\sum_{i\in\{n,p\}}e^{\cos(\Phi_{\text{image}}(I_{t}),E_{i})}}$ (8)
where $\Phi^{l}$ is $l^{th}$ layer of CLIP image encoder, and $a_{l}$ is its corresponding weight.
In the second stage of prompt refinement and enhancement tuning, where learned vectors $E_{p}$ and $E_{n}$ are fine-tuned using margin ranking loss, the formula for $S(I)$ changes as follows:
$S(I)=\frac{e^{\cos(\Phi_{\text{image}}(I),E_{n})}}{\sum_{i\in\{n,p\}}e^{\cos(\Phi_{\text{image}}(I),E_{i})}}$ (9)
In all other aspects the training algorithm is similar to that of CLIP-LIT.
CLIP-LIT-Latent has several advantages over CLIP-LIT. First, CLIP-LIT-Latent does not use the text encoder, which speeds up the training and inference processes, as it is no longer necessary to pass vectors and gradients through the text encoder. This also potentially enables the use of other vision models which do not have text encoder parts rather than CLIP for guidance. Second, images obtained from CLIP-LIT-Latent tend to have more contrast and have better visual quality (see Fig. 4 and Fig. 1). We speculate this results from tuning vectors $E_{p}$ and $E_{n}$ directly in the CLIP latent space. In contrast, in the original CLIP-LIT, vectors $T_{p}$ and $T_{n}$ belong to the text embedding space, and must be tuned so that their projection through the CLIP text encoder satisfies desired conditions. Such optimization is more complex and makes tuning of $T_{p}$ and $T_{n}$ harder and less accurate.
3.3 RAVE: Residual Vector Embedding
(a) Calculation of residual vector $\mathbf{v}_{\textit{residual}}$ based on paired or unpaired training data
(b) Training of enhancement model using guidance from $\mathbf{v}_{\textit{residual}}$
Figure 3: Overview of the RAVE model. (a) First, we calculate residual vector $\mathbf{v}_{\text{residual}}$ based on backlit and well-lit training data. (b) Then we switch to enhancement model training based on the identity loss and the loss based on the residual vector.
In this section we present our model RAVE (ResiduAl Vector Embedding). RAVE includes a further modification of the CLIP guidance. We show that RAVE does not require iterative stages of prompt and model updates. Instead it only requires one stage of enhancement model training.
In RAVE we exploit arithmetic defined in the CLIP latent space. Using well-lit and backlit training data, we construct a residual vector $\mathbf{v}_{\text{residual}}$, which will then be used for enhancement model guidance. $\mathbf{v}_{\text{residual}}$ is defined as follows:
$\begin{split}&\mathbf{v}_{\text{residual}}=f_{\text{norm}}(\mathbf{v}_{\text{well-lit}}-\mathbf{v}_{\text{backlit}})\\ &\mathbf{v}_{\text{well-lit}}=f_{norm}\left(\frac{\sum_{i=1}^{N_{p}}f_{\text{norm}}(\Phi_{\text{image}}(I^{p}_{i}))}{N_{p}}\right)\\ &\mathbf{v}_{\text{backlit}}=f_{norm}\left(\frac{\sum_{i=1}^{N_{n}}f_{\text{norm}}(\Phi_{\text{image}}(I^{n}_{i}))}{N_{n}}\right)\end{split}$ (10)
where $f_{\text{norm}}$ is a normalization function: $f_{\text{norm}}(\mathbf{v})=\frac{\mathbf{v}}{\|\mathbf{v}\|_{2}}$, $N_{p}$ is the number of training well-lit images, $N_{n}$ is the number of training backlit images, $I^{p}_{i}$ is the $i^{th}$ well-lit training image, $I^{n}_{i}$ is the $i^{th}$ backlit training image.
$\mathbf{v}_{\text{residual}}$ is a vector that points in a direction moving from backlit images to well-lit images in the CLIP embedding space. We then use this vector as guidance for the image enhancement model during training. This will train the image enhancement model to produce images with CLIP latent vectors that are close to the CLIP latent vectors of well-lit training images. Overview of the RAVE approach is presented on Fig. 3.
We use the same Unet[21] as enhancement model as in CLIP-LIT and CLIP-LIT-Latent. The enhancement model is trained using the following loss function:
$L_{\text{rave}}=L_{\text{identity}}+\omega L_{\text{residual}}$ (11)
$L_{\text{residual}}=||\Phi_{\text{image}}(I)\cdot\mathbf{v}_{\text{residual}}-\mathbf{v}_{\text{well-lit}}\cdot\mathbf{v}_{\text{residual}}||^{2}$ (12)
where $I$ is an input image, and $v_{\text{well-lit}}$ is a normalized mean of CLIP embeddings of well-lit training images:
$\mathbf{v}_{\text{well-lit}}=f_{\text{norm}}\left(\frac{\sum_{i=1}^{N_{p}}f_{\text{norm}}(\Phi_{\text{image}}(I^{p}_{i}))}{N_{p}}\right)$ (13)
Note that RAVE does not require several repeating stages during training. Before training, the residual vector $\mathbf{v}_{\text{residual}}$ is calculated, and afterwards the training of the image enhancement model begins using $L_{\textit{rave}}$ as the loss function. This makes RAVE training much more efficient than that of CLIP-LIT and CLIP-LIT-Latent. First, the training of the image enhancement model no longer depends on the quality of learned prompt embeddings, as a fixed residual vector is used for guidance. Also, the iterative manner of CLIP-LIT and CLIP-LIT-Latent training regimes makes training less stable, as the image enhancement model training depends on the quality of prompts obtained on the previous step and vice versa. In supplementary material we show that results obtained on intermediate steps of CLIP-LIT training can be of poor quality and can fluctuate from producing under-exposed to over-exposed resulting images from step to step. In contrast, RAVE has a well-defined fixed initial objective. Results show that RAVE converges up to 25 times faster than CLIP-LIT and CLIP-LIT-Latent. Furthermore, images obtained by RAVE have fewer over-exposed areas, producing high quality results.
4 Experiments
Datasets Following CLIP-LIT, we use the BAID[15] and DIV2K[1] datasets for training. We train our models in two settings: using paired and unpaired training data. For the paired setting, we use 380 backlit and corresponding well-lit images from BAID dataset. For the unpaied setting, we use 380 backlit images from the BAID and 384 images from DIV2K dataset. For the unpaired setting, we use the same data as used in CLIP-LIT. For testing, we follow recent literature and use the BAID[15] test dataset, which has 368 backlit and corresponding well-lit images. The data is publicly available.
Training and inference We follow parameter settings of the original CLIP-LIT approach when setting parameters for CLIP-LIT-Latent and RAVE (where applicable), as well as for reproducing CLIP-LIT. Specific details of training and values of hyperparameters are presented in supplementary material.
Metrics Following recent works in image enhancement, we report four metrics: PSNR[9], SSIM[23], and LPIPS[27] (AlexNet version) and FID[16].
Methods for comparison The main model which we compare our approaches to is CLIP-LIT[14], as it showed superior performance on all mentioned metrics, except FID, to many prior image enhancement models. We also include recently published methods such as Restormer[26], Retinexformer[3], DiffIR[24] and Diff-Retinex[25] to the comparison.
4.1 Results
Quantitative comparison The quantitative comparison on the BAID test dataset is presented in Tab. 1. We report test set metrics of the best model checkpoint obtained during training for all the models. Detailed metrics values for all the checkpoints after each training epoch can be found in supplementary material. Note that many approaches can only be trained using paired data, that is why their results are only present in the “paired” section. We see that our method achieves state-of-the-art performance in both settings. In supplementary material, we also present results on time needed to train all the models, and confirm that both CLIP-LIT-Latent and RAVE are much more efficient. Note that in the setting of unpaired data we have two versions of RAVE, referred as RAVE and RAVE shifted in Tab. 1. RAVE shifted is a variant of RAVE with a simple modification of residual vector, see “Interpretation of the residual vector” section below for more details.
Table 1: Quantitative comparison of different methods on the BAID test dataset. The best and second best performances in both settings are in bold and underlined.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
Original images 16.91 0.76 0.21 52.47
Paired Restormer 21.07 0.832 0.192 41.17
Retinexformer 22.03 0.862 0.173 45.27
DiffIR 21.10 0.835 0.175 40.35
Diff-Retinex 22.07 0.861 0.160 38.07
CLIP-LIT 21.93 0.875 0.156 41.16
CLIP-LIT-Latent (ours) 21.84 0.877 0.155 42.56
RAVE (ours) 22.26 0.880 0.139 36.01
Unpaired EnlightenGAN 17.55 0.864 0.196 43.50
CLIP-LIT 21.59 0.874 0.160 46.49
CLIP-LIT-Latent (ours) 21.45 0.877 0.156 46.20
RAVE (ours) 20.39 0.861 0.155 40.11
RAVE shifted (ours) 21.37 0.877 0.149 40.90
(a) Visual comparison of results by CLIP-LIT, CLIP-LIT-Latent and RAVE, all trained with paired data. RAVE produces results with more general contrast.
(b) Visual comparison of results obtained by CLIP-LIT, CLIP-LIT-Latent and RAVE (shifted version), all trained with unpaired data. RAVE produces results with more general contrast.
Figure 4: Visual comparison of results by CLIP-LIT, CLIP-LIT-Latent and RAVE.
Visual comparison For visual comparison, we present representative examples from the BAID test dataset. First, in Fig. 4, we compare CLIP-LIT-Latent and RAVE to original CLIP-LIT. First, note that CLIP-LIT-Latent produces more contrast in general, giving more diverse illumination to different regions. Second, in Fig. 1, we present comparison of images produced by RAVE to images produced by CLIP-LIT. We show that images produced by RAVE have fewer artifacts. We hypothesize that this advantage of RAVE may result from its guiding mechanism: in RAVE we directly point the enhancement model to produce images which have similar CLIP embeddings to that of real well-lit images. This does not allow model to produce regions with artifacts, as its CLIP embeddings won’t match those of real well-lit images. In CLIP-LIT or CLIP-LIT-Latent, on the other hand, the objective of the enhancement model is to produce images for which the CLIP embeddings are simply closer to the positive prompt/vector than to the negative one. This allows the enhancement model to produce less accurate images, CLIP embeddings of which are just “close enough” to the CLIP embeddings of real well-lit images. More results of visual comparison of all the models can be found in supplementary material.
Interpretation of the residual vector In RAVE, we build the residual vector $\mathbf{v}_{residual}$ following formula 16. We noted that $\mathbf{v}_{residual}$ is a vector that points to a direction of moving from backlit images to well-lit images in the CLIP embedding space. Here we prove this by interpreting vector $\mathbf{v}_{residual}$ using the fact that CLIP has a common vector space for embeddings of images and text.
We go through the vocabulary of the CLIP text encoder and find those vocabulary tokens which have the closest and farthest CLIP embeddings to the $\mathbf{v}_{residual}$ in the CLIP latent space in terms of cosine similarity. More formally, we use the following metric:
$\begin{split}sim(token)=f_{\text{norm}}(\Phi_{\text{text}}(e_{\text{token}}))\cdot\mathbf{v}_{\text{residual}}\end{split}$ (14)
Here $e_{token}$ is embedding of vocabulary $token$, which is input to the CLIP text encoder $\Phi_{text}$. We sort vocabulary tokens using $sim$. Table 2 shows 7 vocabulary tokens which have highest and lowest cosine similarity to the $v_{residual}$ vectors obtained from the paired dataset and from the unpaired dataset.
In the case of paired data, many of the least similar tokens have a meaning around "darkness": "silhouette", "dark", "shadow", etc., and absolute value of their cosine similarity to $v_{\text{residual}}$ is around $0.12-0.15$. At the same time, most similar tokens do not share any meaningful semantics, and the absolute value of their cosine similarity to $v_{\text{residual}}$ is much smaller, being around $0.01-0.04$. This supports the hypothesis that $v_{\text{residual}}$ has a meaning of shifting from "dark", "shadowy" images to "normal" images, having average level of lightness. $v_{\text{residual}}$ does not carry any other semantic meaning, as it was calculated using paired dataset, where for each backlit image there is its well-lit analogue, and all the semantics present in the BAID dataset is cancelled as a result of subtraction of the mean well-lit and backlit vectors when constructing $v_{\text{residual}}$. On the other hand, in the case of unpaired data, we see that the least similar tokens have a meaning around "asian": "busan", "beijing", etc., and the most similar ones have a meaning around "nature": "wildlife", "southafrica", "countrylife", etc. Herein, an absolute value of cosine similarities for both sets of tokens if quite high, being around $0.08-0.12$. This reflects the bias in the unpaired datasets. Indeed, in the BAID dataset there are mostly images of Chinese cities or people, while in DIV2K dataset there are many pictures of nature. Due to the bias, RAVE training with unpaired data becomes less effective, making the enhancement model shift images not in an ideal direction in the CLIP latent space. Nevertheless, RAVE still shows good performance when trained with unpaired data.
We also propose a simple modification for correcting the residual vector in the unpaired setting. We simply take $n$ most and least similar tokens to form an additional residual vector as follows:
$\begin{split}&\mathbf{v}_{\text{add\_residual}}=f_{\text{norm}}(\mathbf{v}_{\text{closest}}-\mathbf{v}_{\text{farthest}})\\ &\mathbf{v}_{\text{closest}}=f_{norm}\left(\frac{\sum_{i=1}^{n}f_{\text{norm}}(\Phi_{\text{text}}(T^{c}_{i}))}{n}\right)\\ &\mathbf{v}_{\text{farthest}}=f_{norm}\left(\frac{\sum_{i=1}^{n}f_{\text{norm}}(\Phi_{\text{text}}(T^{f}_{i}))}{n}\right)\end{split}$ (15)
$T^{c}_{i}$ and $T^{f}_{i}$ are tokens with $i^{th}$ most and least similar embeddings to the $\mathbf{v}_{\text{residual}}$. And finally we update the $\mathbf{v}_{\text{residual}}$ as follows:
$\mathbf{v}_{\text{residual}}=\mathbf{v}_{\text{residual}}-cos(\mathbf{v}_{\text{residual}},\mathbf{v}_{\text{add\_residual}})\cdot\mathbf{v}_{\text{add\_residual}}$ (16)
Resulting vector $\mathbf{v}_{\text{residual}}$ will keep the meaning of shifting from "dark", images to "normal" images while losing the unwanted semantics. Experimental results show that RAVE trained with such updated $\mathbf{v}_{\text{residual}}$ produces better results (see RAVE shifted in Tab. 1). In supplementary material we present visual comparison of results obtained by original and shifted versions of RAVE, where it can be clearly seen that RAVE shifted produces more well-lit results, while keeping all the advantages of RAVE. There we also present analysis of tokens with the most and least similar embeddings to the updated $\mathbf{v}_{\text{residual}}$. This analysis supports the claim about the meaning of updated $\mathbf{v}_{\text{residual}}$.
Table 2: Vocabulary tokens and their cosine similarities, which have the lowest and highest cosine similarity to the $v_{\text{residual}}$ vectors calculated for paired and unpaired training datasets.
Paired Unpaired
Lowest similarity Highest similarity Lowest similarity Highest similarity
silhouette -0.151 healthtech 0.035 busan -0.108 wildlifewednesday 0.116
darkness -0.149 whitepaper 0.035 beijing -0.103 southafrica 0.103
dark -0.139 theeconomist 0.030 iu -0.100 ecuador 0.095
webcamtoy -0.139 bona 0.029 guangzhou -0.095 patrol 0.095
blackand -0.132 digitaltransformation 0.027 incheon -0.094 countryfile 0.093
vedere -0.131 sua 0.021 shenzhen -0.094 tasmanian 0.093
shadow -0.130 amarketing 0.021 eunhyuk -0.091 womensart 0.091
[…]
Table 3: Mean similarity of CLIP latent vectors of well-lit and back-lit images from BAID and DIV2K datasets to the CLIP latent vectors of vocabulary tokens
Token BAID backlit BAID well-lit DIV2K well-lit
silhouette 0.254 0.227 0.223
dark 0.247 0.222 0.232
darkness 0.242 0.215 0.227
4.2 Limitations
Despite showing great performance in the task of backlit image enhancement, our models have limitation of non-optimal enhancement of the regions where the loss of information occurs, making restoration ambiguous. Our approach has no generative components and cannot regenerate highly-accurate textures when the information is lost. This leads to sub-optimal performance in such tasks as low-light image enhancement, where images are highly under-exposed (see supplementary material). Moreover, our models are not trained to exactly match pixels of resulting images with ground-truth data, so some sub-optimal lightening of pixels might occur, which results in non-ideal visual quality.
However, existing generative-based models also have limitations. Many of them cannot be trained with unpaired data setting, and require much time and memory to be trained and used. They may also hallucinate detail. As a consequence, the resolution of images it can work with is limited.
5 Conclusion
Based on CLIP-LIT method, we have presented two improved CLIP guidance techniques for backlit image enhancement task, named CLIP-LIT-Latent and RAVE. Our methods show similar or superior performance to that of CLIP-LIT in both cases of training with paired or unpaired data, while having lower time and space requirements. Furthermore, latent vector constructed by RAVE is interpretable and can be used to find biases in the training data, as well as potentially correct them. Moreover, our proposed approach has several advantages over the current SOTA in backlit image enhancement, Diff-Retinex. First, unlike Diff-Retinex, our method can also be trained using unpaired data. Second, it is more lightweight, requiring less compute and space for training and inference, suitable for images with larger resolution.
Acknowledgements
Funding for this research was provided by a Google-DeepMind PhD Studentship, and work utilised Queen Mary’s Andrena HPC facility, supported by QMUL Research-IT. We also sincerely thank Laida Kushnareva for assisting with images and text composition in the paper.